Circular No. 7074

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 981226 L. Di Ciolo, G. Celidonio and G. Gandolfi, BeppoSAX Science Oper-

ation Center, Rome; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands), Utrecht; E. Costa, Istituto Astrofisica Spaziale, Rome; and L. Amati, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna, report: “A γ-ray burst has been detected by the BeppoSAX γ-ray burst monitor and the Wide Field Camera No. 1 on Dec. 26.40793 UT. A GRBM quick-look analysis shows a single event lasting ∼ 20 s, with a peak intensity of 110 counts/s in the 40–700-keV energy band. In the WFC the burst reveals a complex structure and a duration of 260 s. The maximum peak flux is about 0.26 Crab (2–28 keV). The position derived by the WFC image is α = 23^{h}29^{m}40^{s}, δ = −23^{o}55′.5 (equinox 2000.0), with an error radius of 6′ that includes uncertainties due to a nonoptimum attitude control configuration. A follow-up observation with BeppoSAX narrow-field instruments is in progress.”

1996 FG_{3}

P. Pravec and L. Šarounová, Ondřejov Observatory; and M. Wolf, Charles University, Prague, report: “Our photometric observations of this Apollo object from Ondřejov on six nights during 1998 Dec. 10.0-18.9 UT revealed a lightcurve that is a coaddition of two components of different periods: one has the period 0.6728±0.0005 day, with amplitude 0.19 mag, and shows two deep minima and little or no variation at phases between them; the other one has the period 0.14973 ± 0.00002 day, with amplitude 0.06 mag, and most of the signal is in the second harmonic. These periods are consistent with the frequencies found by S. Mottola and F. Lahulla (IAUC 7069); their frequencies apparently correspond to the most significant harmonics, which are the second harmonics in both lightcurve components. Based on our observations, we suggest the hypothesis that 1996 FG_{3} is a binary asteroid. Details on the observations and the interpretation are on

http://sunkl.asu.cas.cz/~ ppravec/96fg3.html.”

COMET 52P/HARRINGTON-ABELL Visual magnitude estimates: Sept. 4.05 UT, 12.4 (M. Reszelski, Szamo-

tuly, Poland, 0.25-m reflector); 23.07, 12.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Oct. 14.06, 13.0 (Hornoch); 28.17, 12.1 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); Nov. 11.84, 11.9 (Hornoch); 30.23, 11.1 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); Dec. 20.82, 10.6 (Reszelski); 26.04, 10.9 (Pereira).

1998 December 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden